Mrs. Rivera is filling planter boxes with soil that is 40% sand, and she buys 10 pounds of a commercial potting soil that is 19% sand.She mixes one of her soil with the 10 pounds commercial potting soil. The resulting soil mixture is 26% sand. How many pounds of soil did Mrs. Rivera use?
Answer by ankor@dixie-net.com(22740)

Mrs. Rivera is filling planter boxes with soil that is 40% sand, and she buys 10 pounds of a commercial potting soil that is 19% sand.
She mixes one of her soils with the 10 pounds commercial potting soil.
The resulting soil mixture is 26% sand.
How many pounds of soil did Mrs. Rivera use?
:
let s = amt of soil used
:
Write a percent sand equation
.40s + .19(10) = .26(s+10)
.40s + 1.9 = .26s + 2.6
.40s - .26s = 2.6 - 1.9
.14s = .7
s = .7/.14
s = 5 pounds of soil used
:
:
See if this checks out
.4(5) + .19(10) = .26(5 + 10)
2 + 1.9 = .26(15)
3.9 = 3.9